OpenStudy (anonymous):

Calculate the average translational kinetic energy of a nitrogen molecule at 27 degree celsius?

OpenStudy (firejay5):

(a) Kinetic energy is a function of temperature alone: K = (3/2) · kb · T = (3/2) · 1.38065×10⁻²³J/K · 300K